### v1

# Towards Resilient Safety-driven Unlearning for Diffusion Models against Downstream Fine-tuning

Abstract
Text-to-image (T2I) diffusion models have achieved impressive image generation quality and are increasingly fine-tuned for personalized applications. However, these models often inherit unsafe behaviors from toxic pretraining data, raising growing safety concerns. While recent safety-driven unlearning methods have made promising progress in suppressing model toxicity, they are identified to be fragile to downstream fine-tuning, where we reveal that state-of-the-art methods largely fail to retain their effectiveness even when fine-tuned on entirely benign datasets. To mitigate this problem, in this paper, we propose ResAlign, a safety-driven unlearning framework with enhanced resilience against downstream fine-tuning. By modeling downstream fine-tuning as an implicit optimization problem with a Moreau Envelope-based reformulation, ResAlign enables efficient gradient estimation to minimize the recovery of harmful behaviors. Additionally, a meta-learning strategy is proposed to simulate a diverse distribution of fine-tuning scenarios to improve generalization. Extensive experiments across a wide range of datasets, fine-tuning methods, and configurations demonstrate that ResAlign consistently outperforms prior unlearning approaches in retaining safety after downstream fine-tuning while preserving benign generation capability well.
[…]
1 Introduction
Text-to-image (T2I) diffusion models have emerged as a dominant class of generative AI due to their unprecedented ability to synthesize high-quality, diverse, and aesthetically compelling general images from natural language descriptions rombach2022ldm ; podell2024sdxl . Beyond synthesizing general images, there is also a growing interest in customizing pretrained models for personalized generation, e.g., generating images of specific facial identities or artistic styles that are underrepresented in the original training data. This is typically achieved by fine-tuning the pretrained base model on a small reference dataset for a few steps ruiz2023dreambooth ; li2024towards ; wang2025sleepermark . The development of several advanced fine-tuning methods hu2022lora ; ruiz2023dreambooth ; han2023svdiff as well as the rapid proliferation of “fine-tuning-as-a-service” platforms civitai2025 has made personalization widely accessible, fueling a surge in applications such as stylized avatars, fan art, and thematic illustrations, which are becoming increasingly popular especially among younger users wang2023factors ; commonsense2024teens ; ali2023constructing .
Yet alongside the rapid advancements of diffusion models, growing concerns have emerged regarding their potential to generate inappropriate or harmful content (e.g., sexually explicit imagery) gandikota2023erasing ; schramowski2023sld ; li2025rethinking . Due to the large-scale and web-crawled nature of their training datasets, modern T2I models inevitably ingest amounts of harmful materials during pretraining schuhmann2022laion . As a result, these models can reproduce such content either when explicitly prompted or inadvertently triggered. For example, recent studies schramowski2023sld ; qu2023unsafe ; li2024art based on real-world user generations demonstrate that widely-deployed models like Stable Diffusion rombach2022ldm are particularly prone to producing unsafe content, even though many of the prompts that lead to unsafe outputs appear benign and may not be intended to generate harmful results. These vulnerabilities not only allow malicious exploitation through direct or adversarial prompting yang2024sneakyprompt ; zhang2024unlearndiff , but also increase the risk of harmful unintended exposure for ordinary, benign users li2024art , raising serious ethical concerns for real-world deployment. In response to these concerns, a variety of safety-driven unlearning methods gandikota2023erasing ; li2024safegen ; zhang2024defensive have recently been proposed to modify the pretrained models’ parameters in order to suppress their capacity for unsafe generation.
While existing methods have shown encouraging results in reducing model toxicity and the resulting unlearned models are promising to be used as “safe” base models for downstream fine-tuning in practical workflows, one natural yet largely unexplored question is whether the safety of unlearned models remains resilient after downstream fine-tuning. Unfortunately, recent studies pan2024leveraging ; gao2024meta have shown that many existing methods can be easily reversed, where fine-tuning on harmful samples for as few as 20 steps can largely recover a model’s unsafe capability. More strikingly, our empirical results reveal that even when fine-tuned on purely benign data, state-of-the-art unlearning methods can regress, with the model’s harmfulness approaching its pre-unlearning state. In other words, even entirely benign users without any malicious intent or harmful data may inadvertently trigger a recovery of unsafe behaviors, posing unforeseen safety risks in real-world use. These findings suggest that current methods may be significantly more brittle than previously assumed and are largely unprepared to serve as reliably safe base models for downstream fine-tuning, underscoring the urgent need for more resilient approaches that can withstand post-unlearning adaptation.
Towards this end, in this paper, we propose a resilient safety-driven unlearning framework dubbed ResAlign to mitigate the aforementioned problem. The intuition behind our method is that unlearning should not only suppress harmfulness at the current model state, but also explicitly minimize the degree to which harmful behaviors can be regained after (simulated) downstream fine-tuning. While conceptually simple, it is particularly challenging to develop a principled and efficient optimization framework to realize this objective. This is because fine-tuning itself is a multi-step optimization process, making it non-trivial to predict which update direction on the original parameters helps minimize the regained harmfulness after downstream fine-tuning. To address this, we approximate fine-tuning as an implicit optimization problem with a Moreau Envelope formulation moreau1965proximite ; rajeswaran2019meta , which enables efficient gradient estimation via implicit differentiation. Besides, to ensure generalizability against the wide variability in real-world downstream fine-tuning procedures (e.g., different datasets, fine-tuning methods, and hyperparameters), we design a meta-learning approach that simulates a distribution of plausible fine-tuning configurations during training, allowing the model to generalize its resilience across a broad range of downstream adaptation scenarios. We also provide insights from the theoretical perspective to explain the empirical effectiveness of our method.
In conclusion, our main contributions are threefold. (1) We empirically reveal that existing safety-driven unlearning methods largely fail to retain their effectiveness after downstream fine-tuning, even when the data does not contain unsafe content. (2) We propose ResAlign, a resilient safety-driven unlearning framework for T2I diffusion models. By leveraging a Moreau Envelope-based approximation and a meta-learning strategy over diverse adaptation scenarios, ResAlign explicitly accounts for and minimizes post-unlearning degradation due to downstream fine-tuning efficiently with high generalizability. We further provide theoretical insights to help understand the empirical effectiveness of our method. (3) Through extensive experiments, we show that ResAlign consistently outperforms baselines in maintaining safety after fine-tuning, and generalizes well to a wide range of advanced fine-tuning methods, datasets, and hyperparameters. It also preserves both general and personalized generation quality well, and remains effective to multiple diffusion architectures, unsafe concepts, and even when the data is harmfully contaminated or adaptively attacked.
2 Background & Related Work
Text-to-Image (T2I) Diffusion Models. Diffusion models are probabilistic generative models that learn the data distribution by reversing a predefined forward noising process through iterative denoising ho2020ddpm ; rombach2022ldm . Given a clean sample $x$ and a predefined noise schedule $\{\alpha_{t},\sigma_{t}\}_{t=1}^{T}$, the forward diffusion process constructs a noisy version at timestep $t$ as $x_{t}=\alpha_{t}x+\sigma_{t}\epsilon$, where $\epsilon\sim\mathcal{N}(0,I)$. Then, a denoising neural network $\hat{\epsilon}_{\theta}$ parameterized by $\theta$ is trained to predict the added noise $\epsilon$ from $x_{t}$ and the timestep $t$. In the context of text-to-image generation rombach2022ldm , the prediction is further conditioned on a natural language prompt $p$. The model thus learns a conditional denoising function $\hat{\epsilon}_{\theta}(x_{t},p,t)$, and its training objective is to minimize the following denoising score matching loss:
$\mathcal{L}_{\text{denoise}}(\theta,\mathcal{D}_{\text{train}})=\mathbb{E}_{\epsilon,t,(x,p)\sim\mathcal{D}_{\text{train}}}\left[w_{t}\cdot\left\|\hat{\epsilon}_{\theta}(\alpha_{t}x+\sigma_{t}\epsilon,p,t)-\epsilon\right\|_{2}^{2}\right],$ (1)
where $x$ is the ground-truth image, $p$ is the text condition, $t$ is uniformly sampled from $\{1,\dots,T\}$, and $w_{t}$ is a weighting factor that balances noise levels. This objective trains the model to accurately predict how to remove noise from noisy images across different timesteps, allowing it to generate realistic images from pure noise when conditioned on text during inference. Currently, the Stable Diffusion series rombach2022ldm ; podell2024sdxl , which operate the diffusion process in a learned latent space and are pretrained on large-scale datasets, stands as the most widely adopted open-source model for T2I generation. They are also increasingly fine-tuned on downstream datasets to learn and generate concepts that are absent in their pretraining datasets for customized generation purposes ruiz2023dreambooth ; peng2025dreambench ; kumari2023multi .
Mitigating Unsafe Generation in T2I Models. To mitigate unsafe generation in large models, several strategies have been developed recently qu2024provably ; yang2024guardti ; liu2024latent ; gandikota2023erasing , which can be broadly categorized into detection-based and unlearning-based. Detection-based strategies deploy external safety filters (e.g., DNN-based harmful prompt/image detectors) to inspect and block unsafe requests compvis2022sc ; liu2024latent . While effective, it does not inherently “detoxify” the diffusion model, often introducing non-negligible additional computational and memory overhead during inference li2024safegen , and can be easily removed when the model is opensourced fernandez2023stable ; rando2022red . On the other hand, unlearning-based approaches internalize safety constraints by directly modifying model parameters to reduce unsafe generation at its source gandikota2023erasing ; kumari2023ablating ; li2024safegen . From a high level, given a pretrained T2I diffusion model $\theta\in\mathbb{R}^{d}$, existing safety-driven unlearning approaches typically aim to optimize the model with the following objective:
$\theta^{*}\in\arg\min_{\theta\in\Theta}\;\mathcal{L}_{\text{harmful}}(\theta)+\alpha\mathcal{R}(\theta),$ (2)
where $\mathcal{L}_{\text{harmful}}(\theta)$ is a harmful loss such that minimizing it encourages the model to unlearn harmful behaviors, and $\mathcal{R}(\theta)$ is an utility-preserving regularization term that maintains the model’s benign generation capabilities. The hyperparameter $\alpha>0$ is a weighting parameter. Since the seminal work of CA kumari2023ablating and ESD gandikota2023erasing , safety-driven unlearning has rapidly gained traction, with a growing line of subsequent works expanding the design space across multiple dimensions including loss function design kumari2023ablating ; gandikota2023erasing , set of updated parameters li2024safegen ; gandikota2023erasing , optimization strategies gandikota2024unified ; zhang2024defensive ; gong2024rece , etc.
Despite these progresses, a recent pioneering work pan2024leveraging has identified a malicious fine-tuning issue, where unlearned models can quickly regain harmful generative abilities after fine-tuning on a small set of unsafe images, producing diverse outputs that resemble the quality and variety of the original pre-unlearned model. To the best of our knowledge, the most advanced attempt to address this issue is LCFDSD pan2024leveraging , which encourages separation in the latent space between clean and harmful data distributions to increase the difficulty of recovering unsafe behaviors. However, this comes with a notable degradation on the benign utility of the model and only brings slight improvements over existing unlearning methods, as validated in our experiments (Sec. 4.2). To summarize, how to effectively mitigate the post-finetuning resurgence of harmful capabilities in unlearned models remains an important yet largely unaddressed problem and is worth further exploration.
3 Methodology
Motivation & Problem Formulation. Despite recent progress in safety-driven unlearning, our experiments (Sec. 4.2) reveal that many state-of-the-art methods fail to retain their effectiveness after downstream fine-tuning, even when the adaptation data is entirely benign (instead of solely with harmful data). We speculate that this fragility stems from their objective formulation (Eq. (2)): they primarily focus on suppressing harmful behaviors at the current parameter state, without accounting for nearby regions in the parameter space that may be reached through downstream fine-tuning. As a result, the local neighborhood of the unlearned model may still be vulnerable, such that even benign gradient signals during fine-tuning can inadvertently push the model into toxic regions and erode its safety. Motivated by this observation, we aim to learn model parameters that are not only safe in their current state, but also lie in regions of the parameter space where safety is preserved under downstream fine-tuning. In other words, we seek to explicitly penalize the increase in harmfulness loss caused by (simulated) downstream fine-tuning during unlearning. Formally, given a pre-trained T2I diffusion model parameterized by $\theta\in\mathbb{R}^{d}$, we propose the following resilient unlearning objective:
$\theta^{*}\in\arg\min_{\theta\in\Theta}\;\mathcal{L}_{\text{harmful}}(\theta)+\alpha\mathcal{R}(\theta)+\beta\big{[}\mathcal{L}_{\text{harmful}}(\theta_{\text{FT}}^{*})-\mathcal{L}_{\text{harmful}}(\theta)\big{]},$ (3)
where $\theta_{\text{FT}}^{*}=\textsc{Adapt}(\theta,\mathcal{D}_{\text{FT}},\mathcal{C})$ represents the parameters obtained by adapting $\theta$ on dataset $\mathcal{D}_{\text{FT}}$ with configuration $\mathcal{C}$. For example, if $\mathcal{C}$ is the standard gradient descent for $T$ steps with a learning rate of $\eta$, then we have $\theta_{\text{FT}}^{*}=\theta-\eta\sum_{t=0}^{T-1}\nabla_{\theta^{(t)}}\mathcal{L}_{\text{FT}}(\theta^{(t)},\mathcal{D}_{\text{FT}})$, where $\mathcal{L}_{\text{FT}}$ represents the standard diffusion denoising loss and $\theta^{(i)}$ indicates the model parameters at the $i$-th step with $\theta^{(0)}=\theta$. $\alpha$ and $\beta$ are hyperparameters that balance the three terms. In this paper, we follow previous works zheng2024imma ; zhang2024defensive and instantiate $\mathcal{L}_{\text{harmful}}(\theta)$ as the negative denoising score matching loss on a set of inappropriate prompt-image pairs and $\mathcal{R}(\theta)$ as the distillation loss that minimizes the difference of noise prediction between the current model and the original model $\hat{\epsilon}_{\theta_{0}}$ on a set of preservation prompts, i.e., $\mathcal{L}_{\text{harmful}}(\theta)=-\mathbb{E}_{\epsilon,t,(x,p)\sim\mathcal{D}_{\text{harmful}}}[w_{t}\cdot\left\|\hat{\epsilon}_{\theta}(\alpha_{t}x+\sigma_{t}\epsilon,p,t)-\epsilon\right\|_{2}^{2}]$ and $\mathcal{R}(\theta)=\mathbb{E}_{x_{T}\sim\mathcal{N}(0,1),t,p\sim\mathcal{P}_{\text{preserve}}}[w_{t}\cdot\left\|\hat{\epsilon}_{\theta}(x_{t},p,t)-\hat{\epsilon}_{\theta_{0}}(x_{t},p,t)\right\|_{2}^{2}]$. Note that our method is also compatible with other designs of these losses, as long as they offer similar utility.
Efficient Hypergradient Approximation via Implicit Differentiation. To optimize the resilient unlearning objective in Eq. (3), one natural idea is to use gradient-based optimization. Thus, we take gradients with respect to $\theta$ and rearrange terms. This yields the following expression:
$(1-\beta)\nabla_{\theta}\mathcal{L}_{\text{harmful}}(\theta)+\alpha\nabla_{\theta}\mathcal{R}(\theta)+\beta\nabla_{\theta}\mathcal{L}_{\text{harmful}}(\theta^{*}_{\text{FT}}).$ (4)
The first two terms in Eq. (4) are standard and can be easily computed via backpropagation. The difficulty lies in the last term, i.e., the hypergradient $\nabla_{\theta}\mathcal{L}_{\text{harmful}}(\theta^{*}_{\text{FT}})$. Because $\theta^{*}_{\text{FT}}$ is an implicit function of $\theta$, according to the chain rule, we have $\nabla_{\theta}\mathcal{L}_{\text{harmful}}(\theta^{*}_{\text{FT}})=\left(\partial\theta^{*}_{\text{FT}}/\partial\theta\right)^{\top}\cdot\nabla_{\theta^{*}_{\text{FT}}}\mathcal{L}_{\text{harmful}}(\theta^{*}_{\text{FT}})$, which involves the Jacobian ${\partial\theta_{\text{FT}}^{*}}/{\partial\theta}$. Since $\theta^{*}_{\text{FT}}$ is obtained through an optimization procedure (e.g., $T$-step gradient descent), this derivative expands into a product of Jacobians ${\partial\theta_{\text{FT}}^{*}}/{\partial\theta}=\prod_{t=0}^{T-1}\left(I-\eta\nabla^{2}_{\theta^{(t)}}\mathcal{L}_{\text{FT}}(\theta^{(t)},\mathcal{D}_{\text{FT}})\right)$, where $\nabla^{2}_{\theta^{(t)}}\mathcal{L}_{\text{FT}}(\theta^{(t)},\mathcal{D}_{\text{FT}})$ is the Hessian matrix of the fine-tuning loss w.r.t. $\theta^{(t)}$. As such, direct computation requires storing and backpropagating through the entire trajectory and their Hessian matrices, which is computationally and memory-prohibitive.
To address this issue, inspired by recent work in bilevel optimization and adversarial training liu2024moreau ; gao2023moreau ; zhang2024stability ; rajeswaran2019meta ; xiao2024uniformly , we approximate fine-tuning as an implicit optimization over the Moreau Envelope moreau1965proximite of the loss. Specifically, we approximately regard the solution of downstream fine-tuning as the minimizer of the following Moreau Envelope (ME) rajeswaran2019meta , that is,
$\theta_{\text{FT}}^{*}\in\arg\min_{\theta^{\prime}}\;\mathcal{L}_{\text{FT}}(\theta^{\prime},\mathcal{D}_{\text{FT}})+\frac{1}{2\gamma}\|\theta^{\prime}-\theta\|^{2},$ (5)
where $\gamma>0$ is a proximity coefficient rajeswaran2019meta . Intuitively, in the context of diffusion model fine-tuning, we typically start from a strong pretrained base model and only take a few gradient steps. This implicitly restricts the solution $\theta_{\text{FT}}^{*}$ to remain near $\theta$, making it well-approximated by the minimizer of the proximal objective in Eq. (5). As a result, $\theta_{\text{FT}}^{*}$ naturally satisfies the first-order optimality condition, i.e., $\nabla_{\theta^{*}_{\text{FT}}(\theta)}\mathcal{L}_{\text{FT}}(\theta_{\text{FT}}^{*},\mathcal{D}_{\text{FT}})+\frac{1}{\gamma}(\theta_{\text{FT}}^{*}-\theta)=0$ according to the Karush-Kuhn-Tucker theorem kuhn2013nonlinear ; rajeswaran2019meta . Since $\theta_{\text{FT}}^{*}$ is implicitly defined as a function of $\theta$ through this optimality condition, we apply the implicit function theorem to differentiate both sides with respect to $\theta$. Then, by right-multiplying both sides by $\nabla_{\theta_{\text{FT}}^{*}}\mathcal{L}_{\text{harmful}}(\theta_{\text{FT}}^{*})$, we can derive the following equation (details in Appendix A):
$\left(\nabla^{2}_{\theta_{\text{FT}}^{*}}\mathcal{L}_{\text{FT}}(\theta_{\text{FT}}^{*},\mathcal{D}_{\text{FT}})+\frac{1}{\gamma}I\right)\cdot\nabla_{\theta}\mathcal{L}_{\text{harmful}}(\theta^{*}_{\text{FT}})=\frac{1}{\gamma}\nabla_{\theta_{\text{FT}}^{*}}\mathcal{L}_{\text{harmful}}(\theta_{\text{FT}}^{*})$ (6)
Let $A:=\nabla^{2}_{\theta^{*}_{\text{FT}}}\mathcal{L}_{\text{FT}}+\frac{1}{\gamma}I$, $x:=\nabla_{\theta}\mathcal{L}_{\text{harmful}}(\theta_{\text{FT}}^{*})$, and $b:=\frac{1}{\gamma}\nabla_{\theta^{*}_{\text{FT}}}\mathcal{L}_{\text{harmful}}$, we observe from Eq. (6) that the hypergradient $\nabla_{\theta}\mathcal{L}_{\text{harmful}}(\theta^{*}_{\text{FT}})$ is essentially the solution to the linear system $Ax=b$, which can be efficiently solved using the Richardson iteration method richardson1911ix :
$x^{(k+1)}=b-\frac{1}{\gamma}\nabla^{2}_{\theta_{\text{FT}}^{*}}\mathcal{L}_{\text{FT}}(\theta^{*}_{\text{FT}},\mathcal{D}_{\text{FT}})\cdot x^{(k)},$ (7)
with initialization $x^{(0)}=0$. Lastly, we take the final iterate $x^{(K)}$ as the implicit gradient $\nabla_{\theta}\mathcal{L}_{\text{harmful}}(\theta_{\text{FT}}^{*})$ and plug it into Eq. (3), which enables end-to-end gradient-based optimization.
Our approach has the following advantages. First, computing the implicit gradient $\nabla_{\theta}\mathcal{L}_{\text{harmful}}(\theta_{\text{FT}}^{*})$ only depends on the final fine-tuned parameters $\theta_{\text{FT}}^{*}$ and the local Hessian-vector product $\nabla^{2}_{\theta_{\text{FT}}}\mathcal{L}_{\text{FT}}\cdot x$. This product can be efficiently computed as a Hessian-vector product (HVP), which is readily supported in modern autodiff frameworks such as PyTorch (via backward-mode Auto Differentiation), without explicitly forming or inverting any second-order Hessian matrices.
Algorithm 1 GetHyperGrad
1:Input: Base parameters $\theta$, downstream fine-tuned model $\theta^{*}_{\text{FT}}$, loss functions $\mathcal{L}_{\text{FT}},\mathcal{L}_{\text{harmful}}$, proximity coefficient $\gamma$, number of steps $K$
2:Output: Hypergradient $\nabla_{\theta}\mathcal{L}_{\text{harmful}}(\theta_{\text{FT}}^{*})$
3:$x^{(0)}\leftarrow 0$, $b\leftarrow\frac{1}{\gamma}\nabla_{\theta_{\text{FT}}}\mathcal{L}_{\text{harmful}}(\theta_{\text{FT}})$
4:for $k=0$ to $K-1$ do
5:    Compute Hessian-vector product $\nabla^{2}_{\theta_{\text{FT}}}\mathcal{L}_{\text{FT}}(\theta_{\text{FT}})\cdot x^{(k)}$
6:     via iterative solver and reverse-mode autodiff
7:    $x^{(k+1)}\leftarrow b-\frac{1}{\gamma}\cdot\nabla^{2}_{\theta_{\text{FT}}}\mathcal{L}_{\text{FT}}(\theta_{\text{FT}})\cdot x^{(k)}$
8:end for
9:Return $x^{(K)}$
Besides, as the gradient is entirely determined by $\theta_{\text{FT}}^{*}$, there is no need to store the intermediate models during fine-tuning, which significantly reduces memory and computational overhead. Finally, we observe that the Richardson iteration used to approximate the implicit gradient (i.e., Eq. (7)) converges quickly within a few steps (5 in this paper), enabling efficient training for large-scale diffusion models. The algorithm procedure for obtaining the hypergradient is placed in Alg.1.
Cross-configuration Generalization via Meta Learning. So far, we have successfully derived a trajectory-independent hypergradient based on a Moreau envelope approximation, which can be efficiently computed as long as $\theta_{\text{FT}}^{*}$ is available. However, this requires access to the downstream fine-tuning configuration $\mathcal{C}$ and dataset $\mathcal{D}_{\text{FT}}$ during unlearning time, which is typically infeasible. A natural workaround is to use a fixed proxy configuration to simulate downstream fine-tuning during training. Yet, such models carry a risk of overfitting to the fixed simulated configuration liu2024metacloak , which may hinder generalization to other settings, especially those unseen during training.
Algorithm 2 ResAlign
1:Input: Initial model parameters $\theta_{0}$, number of outer loop iterations $I$, number of inner loop iterations $J$, dataset $\mathcal{D}$, distribution over configurations $\pi(\mathcal{C})$, loss functions $\mathcal{L}_{\text{FT}},\mathcal{L}_{\text{harmful}},\mathcal{R}$, learning rate $\eta$, hyperparameters $\alpha,\beta,\gamma,K$
2:Output: Final unlearned model parameters $\theta_{I}$
3:for $i=0$ to $I-1$ do $\triangleright$ Outer Loop
4:    $g_{i}\leftarrow(1-\beta)\nabla_{\theta_{i}}\mathcal{L}_{\text{harmful}}({\theta_{i}})+\alpha\nabla_{\theta_{i}}\mathcal{R}({\theta_{i}})$
5:    for $j=1$ to $J$ do $\triangleright$ Inner Loop
6:         Sample data $\mathcal{D}_{\text{FT}}\sim\mathcal{D}$ and configuration $\mathcal{C}\sim\pi(\mathcal{C})$
7:         $\theta^{*}_{\text{FT}_{i,j}}\leftarrow\textsc{Adapt}(\theta_{i},\mathcal{D}_{\text{FT}},\mathcal{C})$
8:         $g_{i,j}\leftarrow\textsc{GetHyperGrad}(\theta_{i},\theta^{*}_{\text{FT}_{i,j}},\mathcal{L}_{\text{FT}},\mathcal{L}_{\text{harmful}},\gamma,K)$
9:    end for
10:    $g_{i}\leftarrow g_{i}+\frac{\beta}{J}\sum^{J}_{j}g_{i,j}$triangleright Aggregating hypergradients
11:    $\theta_{i+1}\leftarrow\theta_{i}-\eta\cdot g_{i}$
12:end for
13:Return $\theta_{I}$
To this end, we draw inspiration from meta-learning finn2017model ; rajeswaran2019meta ; nichol2018first to improve the cross-configuration generalization of ResAlign. Different from conventional meta-learning that aims to enable fast few-shot adaptation, we aim to train a model whose resilience generalizes across a distribution of downstream fine-tuning datasets and configurations. Specifically, we treat the fine-tuning configuration $\mathcal{C}$ (including loss function, optimizer, learning rate, and number of steps) and data sampling $\mathcal{D}_{\text{FT}}$ as meta-variables. During each inner loop iteration, we sample a batch of data $\mathcal{D}_{\text{FT}}\sim\mathcal{D}$ and configuration ${\mathcal{C}}\sim\pi(\mathcal{C})$ from the pool of these meta-variable choices. Then, we run an adaptation process to obtain the fine-tuned model $\theta^{*}_{\text{FT}}=\textsc{Adapt}(\theta,\mathcal{D}_{\text{FT}},\mathcal{C})$ and use Alg. 1 to compute a hypergradient with respect to $\theta$. This process is repeated for $J$ iterations, which gives us a pool of hypergradients reflecting the model’s unlearning dynamics under diverse fine-tuning regimes. Then, these hypergradients are aggregated and used to update the base model parameters in our outer loop iterations. Following previous works finn2017model ; liu2024metacloak , a first-order approximation is used for efficiency. Overall, this meta-generalization mechanism encourages the base model to resist re-acquiring harmful capabilities under a diverse set of plausible fine-tuning conditions while avoiding overfitting to a specific simulation configuration. The overall algorithm procedure is summarized in Alg. 2.
Theoretical Insights. As we will show in experiments, despite introducing only a conceptually simple term, our ResAlign consistently improves safety resilience across a wide range of downstream fine-tuning scenarios. To better understand the (potential) underlying mechanism behind ResAlign’s empirical effectiveness, in this section, we provide a qualitative theoretical analysis, as follows:
Proposition 1.
Let $\theta\in\mathbb{R}^{d}$ and $\theta_{\text{FT}}^{*}\in\mathbb{R}^{d}$ denote the parameters of the base model and the fine-tuned model, respectively. Assume the harmful loss $\mathcal{L}_{\text{harmful}}$ is twice differentiable around $\theta$, the parameter difference of the two models $\xi\in\mathbb{R}^{d}$ is small and can be regarded as a scaled version of a unit random variable $z\in\mathbb{R}^{d}$ with scaling factor $\sigma\in\mathbb{R}$, i.e., $\xi=\sigma z$. If we exploit the second-order Taylor expansion around model parameters and ignore higher-order terms, we have:
$\arg\min_{\theta}~{}\mathbb{E}\big{[}\mathcal{L}_{\text{harmful}}(\theta_{\text{FT}}^{*})-\mathcal{L}_{\text{harmful}}(\theta)\big{]}\approx\arg\min_{\theta}~{}\text{Tr}(\nabla_{\theta}^{2}\mathcal{L}_{\text{harmful}}(\theta)),$ (8)
where $\text{Tr}(\nabla_{\theta}^{2}\mathcal{L}_{\text{harmful}}(\theta))=\sum_{i=1}^{d}\frac{\partial^{2}\mathcal{L}_{\text{harmful}}(\theta)}{\partial\theta_{i}^{2}}$ is trace of the Hessian matrix of $\mathcal{L}_{\text{harmful}}$ w.r.t. $\theta$.
In general, Proposition 1 shows that our additional term can be regarded as an implicit penalty on the trace of the Hessian of the harmfulness loss with respect to $\theta$. The trace of the Hessian is a well-known indicator of the overall curvature of the loss landscape hutchinson1989stochastic ; foret2021sharpnessaware , where larger values reflect sharper minima that are highly sensitive to parameter perturbations, while smaller values correspond to flatter and more stable regions. Interestingly, prior works in generalization and robustness keskar2016large ; foret2021sharpnessaware have shown that SGD-style optimizers tend to lead models to sharp minima, i.e., regions in the loss landscape with high curvature that are sensitive to parameter perturbations. This connection offers a potential explanation for the fragility of existing unlearning methods: since they do not explicitly regularize the curvature of the harmfulness loss, the resulting models, while appearing safe at their current parameter state, may lie in locally sharp and unstable regions of the loss landscape. In such regions, even ordinary parameter updates, like those induced by benign fine-tuning, can result in disproportionately large shifts in harmful loss, thereby inadvertently reactivating harmful capabilities. In contrast, our ResAlign introduces an implicit penalty on the trace of the Hessian of the harmfulness loss, encouraging convergence to flatter regions of the loss surface. This may help reduce the model’s sensitivity to downstream updates and thus improves its resilience to post-unlearning fine-tuning. We provide a proof of Proposition 1 and further discussion in Appendix A.
4 Experiments
4.1 Experimental Setup
Baselines & Diffusion Models. We evaluate our method against 6 text-to-image diffusion models, including the Stable Diffusion v1.4 rombach2022ldm and 5 state-of-the-art baselines for unsafe concept erasure: ESD gandikota2023erasing , SafeGen li2024safegen , AdvUnlearn zhang2024defensive , and two variants of LCFDSD pan2024leveraging (i.e., LCFDSD-NG and LCFDSD-LT). We directly use their provided checkpoints or run their official code with default hyperparameters to obtain their unlearned models. Note that the baselines are mostly implemented and evaluated based on SD v1.4 and mainly focus on the unsafe concept of sexual, and most of them do not provide extensions to other types of unsafe content or to other model architectures. Thus, we conduct our main experiments under the same setting to ensure a fair comparison.
Fine-tuning Methods & Datasets. Both standard fine-tuning and advanced personalization-tuning methods are considered in our experiment. For standard fine-tuning, we adopt (a filtered subset of) two datasets: DreamBench++ peng2025dreambench and DiffusionDB wang2023diffusiondb , which contain high-quality images of human characters, artistic styles, etc. We use these datasets to serve as a representative for benign fine-tuning. In addition, we follow pan2024leveraging and use the Harmful-Imgs dataset, which consists of a diverse set of sexually explicit prompt-image pairs, to understand the resilience of ResAlign when the fine-tuning dataset (partially) contains inappropriate data. Beyond these, we further evaluate the generalizability of our method under more advanced personalization-tuning methods, including LoRA hu2022lora , DreamBooth ruiz2023dreambooth , SVDiff han2023svdiff , and CustomDiffusion kumari2023multi , with more personalization datasets including Pokémon reach_vb_pokemon_blip_captions , Dog peng2025dreambench , ArtBench liao2022artbench , and VGGFace2-HQ chen2023simswap . Finally, we employ the I2P schramowski2023sld and Unsafe qu2023unsafe datasets for measuring harmful generation capabilities, and the COCO chen2015microsoft dataset for assessing benign generation capability.
Evaluation Metrics. We assess model performance from both safety and generation quality perspectives. For harmful generation capability, we report the Inappropriate Rate (IP) and Unsafe Score (US). IP is computed by generating images using the I2P dataset schramowski2023sld prompts and measuring the proportion of outputs flagged as harmful by either the NudeNet nudenetclassifier or Q16 schramowski2022can detector. Similarly, US is computed on images generated from the Unsafe dataset qu2023unsafe prompts with another pretrained NSFW classifier (i.e., MHSC qu2023unsafe ). A lower IP ($\downarrow$) and US ($\downarrow$) indicate weaker unsafe tendencies, meaning the model poses less risk of misuse and is less likely to expose benign users to inappropriate content li2024art ; schramowski2023sld . Note that, different from Pan et al. pan2024leveraging who evaluate model safety on the full set of unsafe prompts, we follow previous works li2024safegen ; zhang2024defensive and mainly focus on the sexual category in both datasets to ensure a more targeted and fair comparison. For benign generation, we use three standard metrics: FID heusel2017gans , CLIP score radford2021learning , and Aesthetics Score laion_ai_aesthetic_predictor , which are widely used by previous works pan2024leveraging ; zhang2024defensive , to measure the ability of the unlearned models in generating benign concepts. Finally, we also follow previous work han2023svdiff ; ruiz2023dreambooth to use CLIP-I and CLIP-T ruiz2023dreambooth , as well as DINO score (oquab2024dinov, ) to measure the performance of our model in generating personalized concepts after fine-tuning.
Tab. 1: Evaluation across different fine-tuning settings. The results are averaged across 3 independent runs.
Model Harmful Generation Benign Generation
Before Fine-tuning Fine-tuned on DreamBench++ Fine-tuned on DiffusionDB FID ↓ / CLIP Score ↑ / Aesthetics Score ↑
IP ↓ US ↓ IP ↓ US ↓ IP ↓ US ↓
SD v1.4 rombach2022ldm 0.3598 0.1850 16.90 / 31.17 / 6.04
ESD gandikota2023erasing 0.0677 0.0100 0.1661 0.0467 0.2209 0.0817 16.88 / 30.26 / 6.01
SafeGen li2024safegen 0.1199 0.0650 0.3154 0.1166 0.3344 0.1333 17.11 / 31.11 / 5.94
AdvUnlearn zhang2024defensive 0.0183 0.0033 0.1038 0.0317 0.2975 0.1233 18.31 / 29.01 / 5.90
LCFDSD-NG pan2024leveraging 0.0788 0.0150 0.2238 0.0867 0.2474 0.0950 47.21 / 30.09 / 5.23
LCFDSD-LT pan2024leveraging 0.1833 0.0467 0.2467 0.0917 0.2832 0.1117 31.69 / 30.72 / 5.60
Ours 0.0014 0.0033 0.0186 0.0050 0.0687 0.0550 18.18 / 31.03 / 5.98
Implementation Details. By default, we initialize our model using SD v1.4 and train it following Alg. 2. For our meta-learning, the distribution of configurations $\pi(\mathcal{C})$ include the learning rates of $[1\times 10^{-4},1\times 10^{-5},1\times 10^{-6}]$, the step of $[5,10,20,30]$, the fine-tuning loss (i.e., $\mathcal{L}_{\text{FT}}$) of both standard denoising loss and the prior-preserved denoising loss ruiz2023dreambooth , the algorithm of both full-parameter fine-tuning and LoRA hu2022lora , and the optimizer of both SGD and Adam. For each adaptation simulation, we select a subset of 50 images from the DiffusionDB wang2023diffusiondb dataset to act as the dataset (i.e., $\mathcal{D}_{\text{FT}}$) and a random combination of the above configurations to form $\mathcal{C}$. The outer loop learning rate and steps are set to $2\times 10^{-4}$ and 120, respectively, and the hyperparameters are empirically set as $\alpha=1,\beta=0.8,K=5,\gamma=1$. The full training process is done on a single NVIDIA RTX A100 GPU with a memory consumption of $\sim 56$GB, and the training time is approximately $58$ GPU minutes. In our evaluation, all fine-tuning is full-parameter fine-tuning on the respective dataset for 200 steps with a batch size of $1$ and a learning rate of $1\times 10^{-5}$ by default. To mitigate the randomness during training and evaluation, are experiments are independently repeated 3 times with different random seeds, and we report the averaged results unless otherwise stated.
4.2 Experimental Results
Main Results. We begin by evaluating the resilience of unlearned models under standard fine-tuning.
Fig. 1: Visualization results on harmful generation. Baseline methods largely lose their effectiveness after fine-tuning while our method retains safety. The black blocks are added by authors to avoid disturbing readers.
As shown in Tab. 1, before downstream adaptation, all methods are able to significantly reduce the inherent toxicity of the pre-trained SD v1.4 model, as reflected by their lower IP and US scores. However, after fine-tuning on both benign datasets, all baselines universally encounter notable resurgence of harmful behaviors, even matching the levels of the original unaligned SD v1.4 (e.g., AdvUnlearn on DiffusionDB and SafeGen on both datasets). We also provide some visualization examples showing how different models react to inappropriate prompts before and after fine-tuning in Fig. 1. These facts indicate that existing unlearning techniques offer limited robustness when subjected to downstream adaptation, even when the dataset does not explicitly contain harmful samples. In contrast, our ResAlign consistently outperforms all baselines by a notable margin. Beyond safety, we also compare the impact of unlearning methods on benign content generation, with quantitative results reported in the rightmost column of Tab. 1 and qualitative results in Fig. 3. While some baselines incur noticeable degradation in FID, CLIP score, or aesthetic quality (e.g., LCFDSD), our method introduces minimal performance drop and remains competitive with the original SD v1.4 and most baselines. This indicates that ResAlign not only ensures stronger safety resilience but also preserves the benign generative capability of the model.
We further analyze the learning dynamics of different methods across various fine-tuning steps. As can be seen in Fig. 2, baseline methods show notable fluctuations and a rapid increase in IP with only
Fig. 2: Evaluation across different fine-tuning steps.
a small number of fine-tuning steps, suggesting that their unlearning is brittle in the weight space. ResAlign, on the other hand, maintains a consistently low level of IP even after extensive fine-tuning (e.g., up to 500 steps), with remarkably smaller fluctuations. These results confirm that our method not only effectively suppresses harmful behavior effectively but also endows the model to better withstand downstream re-acquisition of unsafe capabilities.
Tab. 2: Evaluation on different personalization settings. The results are averaged across 3 independent runs.
Setting (Personalization Method + Dataset) Method Harmful Gen. Personalized Generation
IP $\downarrow$ US $\downarrow$ CLIP-T $\uparrow$ CLIP-I $\uparrow$ DINO $\uparrow$
DreamBooth ruiz2023dreambooth + Dog ruiz2023dreambooth SD v1.4 0.3646 0.1884 0.2671 0.9532 0.8763
Ours 0.0021 0.0260 0.2631 0.9466 0.8605
CustomDiffusion kumari2023multi + ArtBench liao2022artbench SD v1.4 0.3416 0.1900 0.2668 0.6024 0.1611
Ours 0.0011 0.0000 0.2702 0.6171 0.1785
LoRA hu2022lora + Pokémon reach_vb_pokemon_blip_captions SD v1.4 0.3269 0.1977 0.3178 0.6544 0.4549
Ours 0.0032 0.0322 0.3175 0.6512 0.4332
SVDiff han2023svdiff + VGGFace2-HQ chen2023simswap SD v1.4 0.3244 0.1567 0.2491 0.6108 0.5672
Ours 0.0050 0.0083 0.2451 0.5826 0.5510
Effectiveness across More Fine-tuning Methods & Datasets. Beyond standard full-parameter fine-tuning, we further evaluate whether ResAlign remains resilient under more advanced personalization tuning methods, and whether it continues to support benign fine-tuning for downstream tasks.
Fig. 3: Visualization results on benign generation. Our unlearned model maintains both general and personalized generation capability similar to the original SD v1.4.
As shown in Tab. 2, our method consistently maintains strong resilience across all four fine-tuning techniques and datasets. Specifically, the IP and US scores remain close to pre-finetuning levels, demonstrating that ResAlign effectively prevents the re-emergence of harmful behavior under these advanced specialized adaptation settings, including those unseen during training (e.g., SVDiff and CustomDiffusion). At the same time, our model achieves comparable, and in some cases even slightly higher (e.g., on ArtBench), CLIP-I, CLIP-T, and DINO scores compared to the original SD v1.4. These quantitative results, as well as the qualitative visualization results in Fig. 3 suggest that our method well-preserved the personalization capability of the unlearned model, i.e., our ResAlign will not interfere with benign downstream tasks.
Effectiveness across More Fine-tuning Configurations. We also assess whether our method remains effective under other fine-tuning hyperparameters. Specifically, we test two commonly varied settings: learning rate and optimizer. We still use the DreamBench++ and DiffusionDB dataset and apply LoRA fine-tuning with different learning rate and optimizer settings. As shown in Tab. 4 and 5, ResAlign continues to perform well across these configurations, maintaining low harmfulness across different settings. This can be attributed to our meta-learning strategy, which exposes the model to a distribution of configurations, thereby improving generalization to a broad range of potential configs.
Effectiveness on Contaminated Data. Lastly, we consider a challenging scenario where the fine-tuning data is either intentionally or inadvertently contaminated with unsafe content. To evaluate this, we mix our datasets with the Harmful-Imgs dataset pan2024leveraging at varying contamination ratios, and use this contaminated data to fine-tune the unlearned models. As shown in Fig. 4, all methods exhibit increasing levels of harmfulness as the proportion of unsafe data rises. Notably, most baselines experience a steep degradation, with their IP approaching that of the original SD v1.4 when 20% or more of the data is unsafe. In contrast, ResAlign maintains significantly lower harmfulness scores across all contamination levels, demonstrating better resilience even in this challenging scenario.
Overall, we argue perfect resilience against harmful data is inherently difficult—if not impossible—as, an adversary can, afterall, treat the recovery of unsafe behavior as a new learning task deng2024sophon , which pre-trained diffusion models are adept at due to their strong generalization and few-shot
Fig. 4: Evaluation on contaminated data.
learning capability ruiz2023dreambooth ; kumari2023multi . Despite this, we believe the mitigation and insights provided by ResAlign are non-trivial and meaningful. It not only achieves notably better suppression of harmfulness recovery under benign fine-tuning, which is dominant in real-world use cases, but also raises the cost of adversaries by making the re-acquisition of unsafe behaviors harder. We hope our work motivates further investigation into this underexplored yet impactful regime and inspires future strategies that enhance robustness against regaining unlearned harmful capabilities during downstream adaptation.
Tab. 3: Effect of components.
Component Metric
Hyper. Meta. (D) Meta. (C) IP $\downarrow$ FID $\downarrow$
-- -- -- 0.2266 18.24
$\checkmark$ -- -- 0.1826 18.07
$\checkmark$ $\checkmark$ -- 0.0322 18.35
$\checkmark$ $\checkmark$ $\checkmark$ 0.0186 18.18
Fig. 5: Effect of $\gamma$.
Ablation Study. We further conduct an ablation study and hyperparameter analysis on the DreamBench++ dataset. From Tab. 4.2, we can see that every component contributes to our ResAlign, with hypergradient approximation (Hyper.) enabling reducing regained harmfulness and meta-learning on dataset (Meta. (D)) and configurations (Meta. (C)) notably mitigate overfitting and enhance generalization. Besides, the hyperparameter $\gamma$ controls the degree of proximity to the base model. A larger $\gamma$ can better approximate the ground-truth fine-tuned model, yet also leads to higher instability due to larger variance in hypergradient approximation, as shown in Fig. 5. Overall, our method is stable when $\gamma$ is in a reasonable range (i.e., $0\sim 1$). Thus, a wide range of $\gamma$ can be selected for ResAlign.
Tab. 4: Effectiveness of ResAlign under various fine-tuning learning rates. Adapt. refers to an adaptive learning rate scheduler that initiates at $5\times 10^{-3}$ and gradually decays using cosine annealing down to the minimum learning rate of $1\times 10^{-6}$. The reported metric is IP after fine-tuning.
Dataset $1\times 10^{-3}$ $5\times 10^{-4}$ $1\times 10^{-4}$ $5\times 10^{-5}$ $1\times 10^{-5}$ $5\times 10^{-6}$ Adapt.
DreamBench++ 0.018 0.002 0.004 0.004 0.001 0.002 0.011
DiffusionDB 0.059 0.053 0.043 0.054 0.069 0.044 0.059
Tab. 5: Effectiveness of ResAlign under various fine-tuning optimizers. The metric is IP after fine-tuning.
Optimizer DB++ Diff.DB
SGD 0.0004 0.0118
Adam 0.0014 0.0687
AdamW 0.0026 0.0478
5 Conclusion
This paper proposes ResAlign, a novel unlearning framework designed to enhance the resilience of unlearned diffusion models against downstream fine-tuning. Through a Moreau Envelope-based reformulation and a meta-learning strategy, ResAlign effectively suppresses the recovery of harmful behaviors while maintaining benign generation capabilities. The effectiveness of ResAlign is validated through extensive experiments across various datasets and fine-tuning setups, and also qualitatively explained from a theoretical perspective. We hope our work can raise community attention to the resilience of unlearning methods and encourage further research into more robust strategies.
Ethics Statement. This work explores the safety risks of safety-driven unlearned text-to-image diffusion models, particularly their susceptibility of regaining unsafe behaviors after downstream fine-tuning. Our main contribution is ResAlign, a novel framework that enhances the resilience of safety-driven unlearning, mitigating the reemergence of harmful outputs after downstream fine-tuning, which is a defense mechanism. We are confident that our research complies with ethical standards. All experiments are conducted on publicly available datasets, and potentially harmful visualizations are intentionally blurred or blocked. Our method is developed solely to promote the safe and responsible use of generative models and is not intended for any malicious or unethical applications.
Appendix A Omitted Proofs & Derivations
A.1 Omitted Derivation for Eq. (6)
Recall that we approximate the parameters obtained via finite-step fine-tuning from a pretrained model using the following Moreau Envelope (ME) formulation:
$\theta_{\text{FT}}^{*}\in\arg\min_{\theta^{\prime}}\;\mathcal{L}_{\text{FT}}(\theta^{\prime},\mathcal{D}_{\text{FT}})+\frac{1}{2\gamma}\|\theta^{\prime}-\theta\|^{2}$ (9)
As $\theta^{*}_{\text{FT}}$ is the minimizer of the above optimization problem, it satisfies the following first-order optimality condition according to the KKT theorem:
$\nabla_{\theta^{\prime}}\mathcal{L}_{\text{FT}}(\theta_{\text{FT}}^{*},\mathcal{D}_{\text{FT}})+\frac{1}{\gamma}(\theta_{\text{FT}}^{*}-\theta)=0$ (10)
Define a function $F(\theta^{\prime},\theta)=\nabla_{\theta^{\prime}}\mathcal{L}_{\text{FT}}(\theta^{\prime},\mathcal{D}_{\text{FT}})+\frac{1}{\gamma}(\theta^{\prime}-\theta)$. The optimality condition implies that $F(\theta_{\text{FT}}^{*},\theta)=0$. Since $\theta_{\text{FT}}^{*}$ is implicitly defined as a function of $\theta$ through this equation, we can apply the Implicit Function Theorem (IFT) to compute how $\theta_{\text{FT}}^{*}$ changes with respect to $\theta$. Specifically, IFT states that if $F(\theta^{\prime},\theta)=0$ and $\nabla_{\theta^{\prime}}F$ is invertible at $\theta_{\text{FT}}^{*}$, then:
$\frac{\partial\theta_{\text{FT}}^{*}}{\partial\theta}=-\left(\frac{\partial F}{\partial\theta^{\prime}}\right)^{-1}\cdot\frac{\partial F}{\partial\theta}$ (11)
We compute the two Jacobians of $F$ as follows:
$\frac{\partial F}{\partial\theta^{\prime}}=\nabla^{2}_{\theta^{\prime}}\mathcal{L}_{\text{FT}}(\theta_{\text{FT}}^{*},\mathcal{D}_{\text{FT}})+\frac{1}{\gamma}I,\quad\frac{\partial F}{\partial\theta}=-\frac{1}{\gamma}I$ (12)
Substituting into Eq. (11), we obtain:
$\frac{\partial\theta_{\text{FT}}^{*}}{\partial\theta}=\left(\nabla^{2}_{\theta^{\prime}}\mathcal{L}_{\text{FT}}(\theta_{\text{FT}}^{*},\mathcal{D}_{\text{FT}})+\frac{1}{\gamma}I\right)^{-1}\cdot\frac{1}{\gamma}I$ (13)
Now consider the loss function $\mathcal{L}_{\text{harmful}}(\theta_{\text{FT}}^{*})$. By the chain rule:
$\nabla_{\theta}\mathcal{L}_{\text{harmful}}(\theta_{\text{FT}}^{*})=\left(\frac{\partial\theta_{\text{FT}}^{*}}{\partial\theta}\right)^{\top}\cdot\nabla_{\theta_{\text{FT}}^{*}}\mathcal{L}_{\text{harmful}}(\theta_{\text{FT}}^{*})$ (14)
Then, since the Jacobian matrix is symmetric (as it involves the inverse of a symmetric positive definite matrix), we simplify:
$\nabla_{\theta}\mathcal{L}_{\text{harmful}}(\theta_{\text{FT}}^{*})=\left(\nabla^{2}_{\theta^{\prime}}\mathcal{L}_{\text{FT}}(\theta_{\text{FT}}^{*},\mathcal{D}_{\text{FT}})+\frac{1}{\gamma}I\right)^{-1}\cdot\frac{1}{\gamma}\cdot\nabla_{\theta_{\text{FT}}^{*}}\mathcal{L}_{\text{harmful}}(\theta_{\text{FT}}^{*})$ (15)
Finally, multiplying both sides by the inverse term’s denominator yields:
$\left(\nabla^{2}_{\theta_{\text{FT}}^{*}}\mathcal{L}_{\text{FT}}(\theta_{\text{FT}}^{*},\mathcal{D}_{\text{FT}})+\frac{1}{\gamma}I\right)\cdot\nabla_{\theta}\mathcal{L}_{\text{harmful}}(\theta_{\text{FT}}^{*})=\frac{1}{\gamma}\nabla_{\theta_{\text{FT}}^{*}}\mathcal{L}_{\text{harmful}}(\theta_{\text{FT}}^{*})$ (16)
Noting that $\theta^{\prime}$ is evaluated at its optimal value $\theta_{\text{FT}}^{*}$, we equivalently write $\nabla^{2}_{\theta^{\prime}}\mathcal{L}_{\text{FT}}(\theta_{\text{FT}}^{*},\mathcal{D}_{\text{FT}})$ as $\nabla^{2}_{\theta_{\text{FT}}^{*}}\mathcal{L}_{\text{FT}}(\theta_{\text{FT}}^{*},\mathcal{D}_{\text{FT}})$ for clarity. This concludes the derivation. ∎
A derivation similar to ours was presented by Rajeswaran et al. rajeswaran2019meta in the context of meta-learning. Specifically, they model the lower-level transfer (or adaptation) step in meta-learning as a proximal regularized optimization process, which is conceptually similar to our Moreau Envelope-based formulation. This allows them to compute the meta-gradients with respect to the meta-parameters through implicit differentiation, a technique that shares structural similarity with our derivation above. However, while the underlying mathematical tools overlap, our goals and problem settings are fundamentally distinct. The primary objective in Rajeswaran et al. rajeswaran2019meta is to strengthen task-level adaptability in meta-learned models. In contrast, our method is developed to enhance resilience against the reemergence of unsafe behaviors during downstream fine-tuning of generative models.
A.2 Omitted Proof for Proposition 1
Proof.
Prior works have shown that fine-tuning typically induces only small parameter changes to the pretrained parameters radiya2020fine ; fu2023stability ; bai2020taylorized ; tang2023parameter ; fu2023effectiveness . As a result, the fine-tuned parameters can be represented as a sufficiently small additive perturbation of the pretrained ones malladi2023kernel ; fu2023effectiveness ; radiya2020fine , i.e., $\theta_{\text{FT}}^{*}=\theta+\xi$, where $\xi=\sigma z$ [^1] Here, unlike many previous works that restrictively assume $z$ to follow a specific distribution (e.g., Gaussian fort2019deep ; izmailov2018averaging or uniform on the sphere dinh2017sharp ; andriushchenko2020understanding ), or confining analysis to first-order approximations like the Neural Tangent Kernel (NTK) lee2019wide ; ortiz2023task ; jacot2018neural , we only assume the additive perturbation is a scaled unit random variable, i.e., any distribution satisfying $\mathbb{E}[z]=0$ and $\mathbb{E}[zz^{\top}]=\frac{1}{d}I$, which is a milder and more flexible assumption.. By the second-order Taylor expansion around $\theta$ and dropping higher-order terms:
$\mathcal{L}_{\text{harmful}}(\theta_{\text{FT}}^{*})$ $=\mathcal{L}_{\text{harmful}}(\theta+\xi)$
$\approx\mathcal{L}_{\text{harmful}}(\theta)+\nabla_{\theta}\mathcal{L}_{\text{harmful}}(\theta)^{\top}\xi+\frac{1}{2}\xi^{\top}\nabla_{\theta}^{2}\mathcal{L}_{\text{harmful}}(\theta)\xi.$ (17)
Because $z$ is a unit random variable, we have $\mathbb{E}[z]=0$. Then, taking expectation over $z$, we obtain:
$\mathbb{E}\left[\mathcal{L}_{\text{harmful}}(\theta_{\text{FT}}^{*})-\mathcal{L}_{\text{harmful}}(\theta)\right]$ $\approx\sigma\nabla_{\theta}\mathcal{L}_{\text{harmful}}(\theta)^{\top}\mathbb{E}[z]+\frac{\sigma^{2}}{2}\mathbb{E}\left[z^{\top}\nabla_{\theta}^{2}\mathcal{L}_{\text{harmful}}(\theta)z\right]$
$=\frac{\sigma^{2}}{2}\cdot\mathbb{E}\left[z^{\top}\nabla_{\theta}^{2}\mathcal{L}_{\text{harmful}}(\theta)z\right].$ (18)
To compute this expectation, we use the fact that for any symmetric matrix $H$ and random vector $z$ with $\mathbb{E}[zz^{\top}]=\frac{1}{d}I$, we have:
$\mathbb{E}[z^{\top}Hz]=\text{Tr}(\mathbb{E}[zz^{\top}]H)=\text{Tr}\left(\frac{1}{d}I\cdot H\right)=\frac{1}{d}\text{Tr}(H).$ (19)
Applying this to $H=\nabla_{\theta}^{2}\mathcal{L}_{\text{harmful}}(\theta)$, we obtain:
$\mathbb{E}\left[z^{\top}\nabla_{\theta}^{2}\mathcal{L}_{\text{harmful}}(\theta)z\right]=\frac{1}{d}\text{Tr}\left(\nabla_{\theta}^{2}\mathcal{L}_{\text{harmful}}(\theta)\right).$ (20)
Hence, substituting into Eq. (A.2), we obtain the following expression:
$\mathbb{E}\left[\mathcal{L}_{\text{harmful}}(\theta_{\text{FT}}^{*})-\mathcal{L}_{\text{harmful}}(\theta)\right]\approx\frac{\sigma^{2}}{2d}\cdot\text{Tr}\left(\nabla_{\theta}^{2}\mathcal{L}_{\text{harmful}}(\theta)\right).$ (21)
Since the term $\frac{\sigma^{2}}{2d}$ is a constant independent of $\theta$, minimizing the expected regained harmful loss is equivalent to minimizing the trace of the Hessian matrix:
$\arg\min_{\theta}~{}\mathbb{E}\left[\mathcal{L}_{\text{harmful}}(\theta_{\text{FT}}^{*})-\mathcal{L}_{\text{harmful}}(\theta)\right]\approx\arg\min_{\theta}~{}\text{Tr}\left(\nabla_{\theta}^{2}\mathcal{L}_{\text{harmful}}(\theta)\right).$ (22)
Thus, we finished the proof. ∎
Fig. 6: Illustration of the impact of (benign) fine-tuning on harmful loss for flat (left) and sharp (right) minima.
Remarks & Discussions. Proposition 1 suggests that our proposed additive term implicitly penalizes the trace of the Hessian of the harmful loss near the unlearned model parameters. In doing so, it encourages the optimization process not only to minimize the harmful loss, but also to favor solutions with flatter local geometry, i.e., regions with lower curvature. This intuition is illustrated in Fig. 6. While prior unlearning methods aim to reduce $\mathcal{L}_{\text{harmful}}$, they typically rely on standard SGD-based optimizers, which are known to prefer sharp local minima (keskar2016large, ). As a result, these methods may converge to parameter regions such as $\theta^{\prime*}_{\text{unlearn}}$ that, although locally optimal, are highly sensitive to downstream perturbations. As a result, even benign fine-tuning, which aims to optimize only the benign loss $\mathcal{L}_{\text{FT}}$ (i.e., aim to pushing $\theta^{\prime*}_{\text{unlearn}}$ to $\theta^{\prime*}_{\text{FT}}$), may inadvertently push the model back into regions of high harmful loss, reactivating undesired behaviors.
In contrast, our ResAlign method implicitly promotes flatter optima (e.g., $\theta^{*}_{\text{unlearn}}$ in the figure). These flatter solutions are more resilient to subsequent fine-tuning: local updates to minimize $\mathcal{L}_{\text{FT}}$ induce smaller changes in $\mathcal{L}_{\text{harmful}}$, thereby reducing the risk of unlearning failure. This perspective suggests that the fragility of prior unlearning methods may arise not from the nature of the downstream data itself, but from the sharpness of the solutions they converge to.
With that being said, we acknowledge that our proposition, like other analyses based on Taylor expansions and additive perturbation models jacot2018neural ; kirkpatrick2017overcoming ; foret2020sharpness , may become less accurate in scenarios involving large parameter shifts (e.g., very large learning rates or very long fine-tuning schedules), or when the perturbation lacks a well-behaved distribution kirkpatrick2017overcoming ; stephan2017stochastic . Nonetheless, in practice, particularly in diffusion model fine-tuning setting, small learning rates and moderate step sizes are commonly used to preserve generative quality and prevent overfitting ruiz2023dreambooth ; han2023svdiff ; hu2022lora , which supports the validity of the small perturbation assumption. Moreover, since our meta-learning process is performed over diverse data sampling and configurations, modeling $\xi$ as an isotropic random variable is a reasonable abstraction, as widely adopted by previous works yoon2018bayesian ; amit2018meta ; rajeswaran2019meta . As one of the earliest attempts to theoretically understand the interplay between unlearning and subsequent fine-tuning, our analysis is meant to provide a qualitative, approximate explanation rather than definitive conclusions. We hope our work can inspire the community to pursue deeper and more precise theoretical investigations into the robustness of unlearning under realistic deployment conditions.
Appendix B Discussion
B.1 Limitations & Future Work
Our work still has the following limitations, which we aim to address in future work. First, our approach requires simulating fine-tuning during the unlearning process, which introduces additional computational overhead. However, the total cost remains within a practical range ($\sim 56$ GPU minutes for a full run in our main experiments), which is significantly more efficient than retraining a model from scratch. In real-world scenarios, users can choose methods that best balance their needs for resilience and safety given their computational budget. Second, our Moreau envelope-based approximation is inherently an approximation. Nevertheless, our experiments show that it is robust to long fine-tuning steps and a wide range of learning rates, and it outperforms methods based on first-order approximations. As discussed in the main paper, obtaining fully accurate hypergradients would require tracking and computing full Hessian matrix products, which is computationally prohibitive for diffusion models. Future work could explore more accurate yet efficient estimation strategies. Third, we acknowledge that when the downstream fine-tuning dataset is entirely composed of toxic data, both our method and existing baselines inevitably experience an increase in harmfulness. We believe that perfect resilience under such adversarial conditions is intrinsically difficult, given the strong generalization and few-shot capabilities of modern diffusion models. Nonetheless, our method demonstrates stronger resilience compared to baselines, maintaining relatively low IP even under fully toxic fine-tuning. Forth, similar to previous works, the NSFW classifiers used in our experiments are not perfectly accurate, and there might be false negatives or false positives. However, as all methods are evaluated under the same setting, the results are still fair and comparable to a large extent. We hope future work can build more accurate classification models. Finally, while we designed and evaluated adversarial & adaptive attacks and demonstrated ResAlign’s robustness against them, we recognize that stronger or more sophisticated attacks may emerge in the future. Developing robust defenses against such potential threats remains an important direction for future work.